Question with a problem I'm having.
just about one aspect.
1/2 * 22 /2 + 1/2 *22 /2 = 110/16
how did the 110 come about?

Answer by checkley77(12844)

1/2 * 22 /2 + 1/2 *22 /2 = 110/16
----------------------------------
1/2*22/2+1/2*22/2
22/4+22/4
44/4=11 ans.